# Exploring the Low-Pass Filtering Behavior in Image Super-Resolution

### v1

@@ 4.2.2 The merits of FSDS
Previous image evaluation metrics, such as PSNR, SSIM [25], have focused on statistical or structural features of images, but no work has evaluated images from the perspective of their frequency spectrum. The spectrum is the distribution of signals or images at different frequencies, focusing on specific frequency components. It is crucial for capturing details, eliminating noise, and comprehensively understanding image features. In image processing, spectrum analysis provides a more accurate evaluation, particularly playing a key role in applications sensitive to details. Due to the nature of Fourier transformation, which involves every pixel of the image in the computation, it encompasses not only information such as signal-to-noise ratio and structural similarity but also the overall similarity of the entire image. Therefore, evaluating image quality from the perspective of the spectrum is highly reasonable and necessary. Our FSDS metric can reflect distribution differences by employing a paradigm of integrating first in the frequency spectrum and then comparing. In other words, FSDS not only reflects the signal-to-noise ratio captured by the PSNR metric and the structural similarity indicated by the SSIM metric, but also captures features that these two metrics cannot represent. In the next paragraph, we will use two toy examples to demonstrate the rationale and advantages of FSDS.
@@ 5 Experiments
Due to page limitations, we can only present three of the most crucial experiments in this section, namely: 1) the relationship between the low-pass filter passband width and ISR performance; 2) various network impulse responses; 3) a comparison of FSDS metrics with PSNR and SSIM metrics on the DIV2K dataset. For more experiments, please refer to Appx. C.
@@ 5.2 Experiment on Impulse Response
We select several mainstream backbones and their derivatives commonly used for ISR tasks [13, 5, 19, 26, 39, 7, 12, 10, 38] and conduct impulse response experiments. The experimental results are compared with the sinc function and depicted in Fig. 9. The input image is an $11\times 1$ 11 image where only the pixel at position $(5,5)$ is white (the values for all three channels at this position are 255, with indices starting from 0), and the rest of the image is black (with values of 0). According to Tab. 3, it can be observed that as the ISR factor increases, the central peak of the output Sinc function becomes wider and more pronounced. To balance visual saliency and the maximum ISR factor achievable by certain networks, we opted for a 4x ISR factor. Observing the experimental results, we can notice that regardless of the neural network structure used for ISR, whether it’s a CNN or a transformer, the impulse response exhibits some degree of similarity to the two-dimensional Sinc function. This similarity is particularly pronounced in networks like RDN [39] and RCAN [38]. Despite some distortion in comparison to the Sinc function, EDSR [13], EQSR [24], and their derivatives still exhibit significant features of the Sinc function, including the central bright spot and elongated bright patches in the cardinal directions. Considering the parameter counts among different backbones (EDSR $<$ SwinIR $<$ RDN), we posit that the clarity of the Sinc function is related to the network’s parameter count – the larger the parameter count, the clearer the Sinc function. Thus, we can summarize that the clearer the sinc patter is, the more it indicates its effectiveness in reconstructing low-frequency information.
@@ 5.3 Experiment on FSDS Metric
Table 1: Comparison of PSNR, SSIM [25], LPIPS [37] and FSDS metrics for different methods on the DIV2K dataset [1]. Items with the highest and the second-highest mean values are highlighted in red and blue respectively. The gray superscripts are the order of each method.
We conducted tests on the validate set of the DIV2K dataset [1] using several methods [13, 5, 19, 26, 10, 39, 12, 32], depicted in Tab. 1. The evaluation metrics include PSNR, SSIM [25], LPIPS [37], and our FSDS. All tests are performed using code and weights available in open-source official repositories. For all methods, we conduct experiment for $\times$2 to $\times$4. For methods that support arbitrary-scale ISR, we test for $\times$6 and $\times$12 as well. In the $\times$2 to $\times$4 range, GRLBase[11] consistently achieves the best performance across PSNR, SSIM and LPIPS metrics, and FSDS shows that SwinIR [12] achieves the best performance. For $\times$6 and $\times$12, RDN-LTE [10] exhibits the best PSNR and SSIM metrics, while RDN-LIIF performs best on LPIPS and the FSDS metric.
From Section 4.2, we claim that previous metrics are not sensitive to high-frequency information, while FSDS does. This can be proven by Tab. 1. In the case of slight high-frequency loss, such as on scales $\times$ 2 to $\times$ 4, FSDS responds differently compared to previous metrics. In cases that suffer from severe high-frequency loss, such as on $\times$ 6 and $\times$ 12 scales, FSDS shows consistency with previous metrics. When high-frequency loss is slight, previous metrics fail to reflect such high-frequency loss and while the loss becomes more severe, they start to capture such loss. This observation shows the necessity of applying the FSDS metrics to assess image quality objectively.
@@ 5.4 Some Exceptions to Impulse Responses
Figure 10: The impulse response of SwinIR-Real network[12] is not a obvious ‘sinc’ function.
We observe that not all impulse response of networks is ‘sinc’ function, as shown in Fig. 10. The SwinIR-Real [12] is trained using adversarial loss while methods in Fig. 9 uses loss like $\ell_{1}$ or $\ell_{2}$ loss.
@@ 6 Conclusion
In this paper, we report an intriguing observation – the sinc phenomenon, revealing that the impulse response of ISR networks act as low-pass filters. Building on this insight, we introduce a novel analytical approach called Hybrid Response Analysis (HyRA). HyRA treats a neural network as a combination of a linear system and a non-linear system with a zero impulse response. The linear system functions as a low-pass filter, while the non-linear system utilizes prior knowledge to inject high-frequency details. To assess the neural network’s information recovery across the frequency spectrum, we propose the Frequency Spectrum Distribution Similarity (FSDS) metric. This metric uncovers properties overlooked by previous metrics, and experiments validate the rationality and necessity of the FSDS metric.
@@ Appendix A Notation Conventions
$sinc_{\omega}(t)$ The sinc funtion with parameter $\omega$, $sinc_{\omega}(t)=\frac{sin(\omega t)}{\pi t}$. The sinc function is the time-domain waveform of a ideal low-pass filter.
@@ Appendix G Comparison Between FSDS and $\ell_{1}$, $\ell_{2}$ norms
We compare our proposed FSDS metric with $\ell_{1}$ norm, and $\ell_{2}$ norm on both frequency domain and image domain as depicted in Tab. 4 and Tab. 5. From these two figures, we can observe that $\ell_{1}$ norm and $\ell_{2}$ norm produces similar ranking orders when they are calculated on the same domain, indicating that $\ell_{1}$ norm is equivalent to $\ell_{2}$ norm when assessing image quality. However, the ranking orders produced by FSDS is distinctive to that of $\ell_{1}$ and $\ell_{2}$. This means our FSDS metric reflects image quality from a unique way.
@@
Appendix H Code Repositories

### v4

@@ 4.2.2 The merits of FSDS
Previous image evaluation metrics, such as PSNR, SSIM [26], have focused on statistical or structural features of images, but no work has evaluated images from the perspective of their frequency spectrum. The spectrum is the concentrated expression of components with different changing rates in a signal or image. It is crucial for capturing details, eliminating noise, and comprehensively understanding image features. In image processing, spectrum analysis provides a more accurate evaluation, particularly playing a key role in applications sensitive to details. Due to the nature of Fourier transformation, which involves every pixel of the image in the computation, it encompasses not only information such as signal-to-noise ratio and structural similarity but also the overall similarity of the entire image. Therefore, evaluating image quality from the perspective of the spectrum is highly reasonable and necessary. Our FSDS metric can reflect distribution differences by employing a paradigm of integrating first in the frequency spectrum and then comparing. In other words, FSDS not only reflects the signal-to-noise ratio captured by the PSNR metric and the structural similarity indicated by the SSIM metric, but also captures features that these two metrics cannot represent. In the next paragraph, we will use two toy examples to demonstrate the rationale and advantages of FSDS.
@@ 5 Experiments
Due to the page limitation, we can only present three of the most crucial experiments in this section, namely: 1) the relationship between the low-pass filter passband width and ISR performance; 2) various network impulse responses; 3) a comparison of FSDS metrics with PSNR, SSIM and LPIPS [36] metrics on the DIV2K dataset. For more experiments, please refer to Appx. C.
@@ 5.2 Experiment on Impulse Response
We select several mainstream backbones and their derivatives commonly used for the ISR task [13, 5, 20, 27, 38, 7, 12, 10, 37] and conduct impulse response tests. The experimental results are compared with the sinc function and depicted in Fig. 9. The input image is an $11\times 1$ 11 image where only the pixel at position $(5,5)$ is white (the values for all three channels at this position are 255, with indices starting from 0), and the rest of the image is black (with values of 0). According to Tab. 3, it can be observed that as the ISR factor increases, the central peak of the output sinc function becomes wider and more pronounced. To balance visual saliency and the maximum ISR factor achievable by certain networks, we opted for a 4x ISR factor. Observing the experimental results, we can notice that regardless of the neural network structure used for ISR, whether it’s a CNN or a transformer, the impulse response exhibits some degree of similarity to the two-dimensional sinc function. This similarity is particularly pronounced in networks like RDN [38] and RCAN [37]. Despite some distortion in comparison to the sinc function, EDSR [13], EQSR [25], and their derivatives still exhibit significant features of the sinc function, including the central bright spot and elongated bright patches in the cardinal directions. From Tab. 1, we observe that networks exhibiting superior performance tend to generate impulse responses that closely resemble the sinc function. This observation suggests that preserving low-frequency information more effectively can also enhance performance. However, few previous works has focus on low-frequency, giving us a new idea for furture ISR networks.
@@ 5.3 Experiment on FSDS Metric
Table 1: Comparison of PSNR, SSIM [26], LPIPS [36] and FSDS metrics for different methods on the DIV2K dataset [1]. Items with the highest and the second-highest mean values are highlighted in red and blue, respectively. The gray superscripts are the order of each method.
We conducted tests on the validation set of the DIV2K dataset [1] using several methods [13, 5, 20, 27, 10, 38, 12, 32], depicted in Tab. 1. The evaluation metrics include PSNR, SSIM [26], LPIPS [36], and our FSDS. All tests are performed using code and weights available in open-source official repositories. For all methods, we conduct experiment for $\times$2 to $\times$4. For methods that support arbitrary-scale ISR, we test for $\times$6 and $\times$12 as well. In the $\times$2 to $\times$4 range, GRLBase[11] consistently achieves the best performance across PSNR, SSIM and LPIPS metrics, and FSDS shows that SwinIR [12] achieves the best performance. For $\times$6 and $\times$12, RDN-LTE [10] exhibits the best PSNR and SSIM metrics, while RDN-LIIF performs best on LPIPS and the FSDS metric.
From Section 4.2, we claim that previous metrics are not sensitive to high-frequency information, while FSDS does. This can be proven by Tab. 1. In the case of slight high-frequency loss, such as on scales $\times$ 2 to $\times$ 4, FSDS responds differently compared to previous metrics. In cases that suffer from severe high-frequency loss, such as on $\times$ 6 and $\times$ 12 scales, FSDS shows consistency with previous metrics. This is because when high-frequency loss is slight, previous metrics fail to reflect such high-frequency loss and while the loss becomes more severe, they start to capture such loss. This observation shows the necessity of applying the FSDS metrics to assess image quality objectively.
@@ 5.4 Some Exceptions to Impulse Responses
Figure 10: The impulse response of SwinIR-Real [12] and ESRGAN [24] is not an obvious sinc function.
We observe that not all impulse response of networks is ‘sinc’ function, as shown in Fig. 10. SwinIR-Real [12] and ESRGAN [24] are trained using adversarial loss, while methods in Fig. 9 uses loss like $\ell_{1}$ or $\ell_{2}$ loss. Therefore, we believe the ‘sinc’ impulse response is related to the loss function.
@@ 6 Conclusion
In this paper, we report an intriguing observation. i.e., the sinc phenomenon, which reveals that the impulse response of ISR networks act as low-pass filters. Building on this observation, we introduce a novel approach called Hybrid Response Analysis (HyRA) to explore the hidden behavior of ISR networks. HyRA treats a neural network as a combination of a linear system and a non-linear system with a zero impulse response. The linear system functions as a low-pass filter, while the non-linear system utilizes prior knowledge to inject high-frequency details. To assess the neural network’s information recovery across the frequency spectrum, we propose the Frequency Spectrum Distribution Similarity (FSDS) metric. FSDS uncovers properties overlooked by previous metrics, and experiments validate the rationality and necessity of it.
@@ Appendix A Notation Conventions
$sinc_{\omega}(t)$ The sinc funtion with parameter $\omega$, $sinc_{\omega}(t)=\frac{sin(\omega t)}{\pi t}$. The sinc function is the time-domain waveform of an ideal low-pass filter.
@@ Appendix G Comparison Between FSDS and $\ell_{1}$, $\ell_{2}$ norms
We compare our proposed FSDS metric with $\ell_{1}$ norm, and $\ell_{2}$ norm on both frequency domain and image domain as depicted in Tab. 4 and Tab. 5. From these two figures, we can observe that $\ell_{1}$ norm and $\ell_{2}$ norm produces similar ranking orders when they are calculated on the same domain, indicating that $\ell_{1}$ norm is equivalent to $\ell_{2}$ norm when assessing image quality. However, the ranking orders produced by FSDS is distinctive to that of $\ell_{1}$ and $\ell_{2}$. This means our FSDS metric reflects image quality in a unique way.
@@
Appendix H How G(I) is Trained and How G(I) and H(I) Vary Dependently
Figure 19: $N(I)$, $H(I)$ and $G(I)$ from different epoches during training.
To better explain how $G(I)$ is trained and how $G(I)$ and $H(I)$ vary dependently, we conduct a new experiment on observing their varying progress during training. We train a vanilla RDN [38] network from scratch and obtain the impulse response, $H(I)$, $G(I)$, and $N(I)$ from each epoch (see Fig. 19). As shown in the figure, in the second row, the $sinc$ phenomenon becomes clearer along with the training process, this indicates that the network is gradually learning the low-pass filter. The same conclusion can be further supported by observing the variation in $H(I)$ as shown in the third row. In this row, $H(I)$ illustrates the phenomenon of the grid-like distortion vanishing while low-frequency areas getting smoother. The fourth row depict the plot of $G(I)$. This row demonstrates that $G(I)$ is capturing more and more high-frequency information, such as edges. This is very significant especially when comparing the result from Epoch 1 and Epoch 50. Such increase in high-frequncy information can also be found in the frequency spectrum. During the entire training process, the network fails to recover the words on the wall (please see the magnified area, there are words on the wall next to the blue awning). Observing $G(I)$, we can find no sign of words as well. This indicates that the network treats it as low-frequency information and ignores it, pointing the way for future network improvements.
Appendix I Code Repositories